OpenStudy (anonymous):

How do i use implicit differentiation to find f' of cos(x) + sqrt(y) = 5

OpenStudy (anonymous):

\[d/dx[\cos(x)+\sqrt{y}]=5d/dx\] \[-\sin(x)+1/2(y^{-1/2})dy/dx=0\] \[1/2(y^{-1/2})dy/dx=\sin(x)\] \[dy/dx=2\sqrt{y} \sin(x)\]

OpenStudy (anonymous):

Ok thats what I thought, thank you!

OpenStudy (anonymous):

no problem